. Find the solution of the linear system using Gauss-Jordan elimination

𝑣 + 2𝑀 βˆ’ 2π‘₯ βˆ’ 𝑦 = 0

2𝑣 + 3𝑀 βˆ’ 5π‘₯ + 𝑦 βˆ’ 7𝑧 = 0

𝑀 + π‘₯ βˆ’ 𝑦 + 𝑧 = 0

βˆ’π‘£ + 𝑀 + 5π‘₯ βˆ’ 𝑦 = 0




1
Expert's answer
2021-04-14T12:51:49-0400

Solution.

"\ud835\udc63 + 2\ud835\udc64 \u2212 2\ud835\udc65 \u2212 \ud835\udc66 = 0"

"2\ud835\udc63 + 3\ud835\udc64 \u2212 5\ud835\udc65 + \ud835\udc66 \u2212 7\ud835\udc67 = 0"

"\ud835\udc64 + \ud835\udc65 \u2212 \ud835\udc66 + \ud835\udc67 = 0"

"\u2212\ud835\udc63 + \ud835\udc64 + 5\ud835\udc65 \u2212 \ud835\udc66 = 0"


"\begin{bmatrix}
 1 & 2&-2&-1&0&0 \\
 2 & 3&-5&1&-7&0\\
 0&1&1&-1&1&0\\
 -1&1&5&-1&0&0
\end{bmatrix}"

Find the pivot in the 1st column in the 1st row

"\begin{bmatrix}
 1 & 2&-2&-1&0&0 \\
 2 & 3&-5&1&-7&0\\
 0&1&1&-1&1&0\\
 -1&1&5&-1&0&0
\end{bmatrix}"

Eliminate the 1st column

"\begin{bmatrix}
 1 & 2&-2&-1&0&0 \\
 0 & -1&-1&3&-7&0\\
 0&1&1&-1&1&0\\
 0&3&3&-2&0&0
\end{bmatrix}"

Find the pivot in the 2nd column in the 2nd row (inversing the sign in the whole row)

"\begin{bmatrix}
 1 & 2&-2&-1&0&0 \\
 0 & 1&1&-3&7&0\\
 0&1&1&-1&1&0\\
 0&3&3&-2&0&0
\end{bmatrix}"

Eliminate the 2nd column

"\begin{bmatrix}
 1 & 0&-4&5&-14&0 \\
 0 & 1&1&-3&7&0\\
 0&0&0&2&-6&0\\
 0&0&0&7&-21&0
\end{bmatrix}"

Make the pivot in the 4th column by dividing the 3rd row by 2

"\begin{bmatrix}
 1 & 0&-4&5&-14&0 \\
 0 & 1&1&-3&7&0\\
 0&0&0&1&-3&0\\
 0&0&0&7&-21&0
\end{bmatrix}"

Eliminate the 4th column

"\begin{bmatrix}
 1 & 0&-4&0&1&0 \\
 0 & 1&1&0&-2&0\\
 0&0&0&1&-3&0\\
 0&0&0&0&0&0
\end{bmatrix}"

Solution set:

"v = 4x - z;\nw = - x + 2z;\ny = 3z;\n\nx, z - free."

Answer: "v = 4x - z;\nw = - x + 2z;\ny = 3z;\n\nx, z - free."